# Towards Reliable Proof Generation with LLMs: A Neuro-Symbolic Approach

### v1

@@ 1 Introduction
Figure 2: An example problem from FormalGeo-7k (level 3). Left: Problem inputs, including a natural language description and formal representations – construction (entities and relations), extended construction (inferred based on extension rules), conditions, and goal. The input also includes a numeric answer (or expression) and a formal proof, composed of steps, each invoking a theorem from the dictionary. Right: The Theorem Dictionary. Theorems include variation id, arguments, premises and conclusions. The full dataset also includes diagrams, but we do not use them here. A recent work Xu et al. (2024) has shown that multimodal LLMs often struggle with the visual aspects of math problems, often performing better without the images.
@@
Mathematical proofs demand rigorous logical deduction, symbolic manipulation, and an understanding of abstract concepts that go beyond statistical correlations in language. The requirement for absolute truth and the absence of ambiguity in mathematical reasoning present a significant challenge for models trained to generate plausible text rather than formally valid inferences Singh et al. (2024); Jiang et al. (2025); Pan et al. (2025). In addition, the often lengthy nature of proofs necessitates a level of sustained logical coherence and hierarchical reasoning that is hard for current LLMs.
Enabling LLMs to generate rigorous and verifiable proofs can dramatically boost LLMs’ reliability, accuracy and consistency, unlocking applications in mathematics, science and education Welleck et al. (2022); Gupta and MacLellan (2025); Hofman et al. (2023), as well as many safety-critical and security-critical tasks.
@@
As a proof of concept, we focus on Euclidean geometry – a domain that is symbolic, verifiable, and rich in structural analogies. Note that our primary objective in this paper is to evaluate whether our symbolic augmentations can improve the proof-generation capabilities of general-purpose LLMs. As such, while geometry is the domain we test our ideas in, our focus is not on competing with state-of-the-art, specialized geometry solvers. Rather, we wish to quantify the gains enabled by our method. Our main contributions are:
@@ 4 Experimental Setup
Base model. We use OpenAI’s o1 model with the default parameters as our base model. o1 is a state-of-the-art LLM known for its capabilities in mathematical reasoning, proof generation, and complex problem-solving OpenAI (2024b). We experimented with several models, including GPT-4, GPT-4o-mini, and GPT-4o OpenAI (2023, 2024a), and observed that these models tend to produce proofs with a significant number of syntax errors. o1 consistently outperformed them, making it the strongest baseline for our task. We leave the evaluation of additional base models for future work.
@@ 5 Results
In setting (1) our method achieved an overall accuracy of 48%, substantially outperforming the baseline’s 10% (blue circle vs. red hollow circle), $p=$3.81e-06 in the McNemar test at the 0.05 level. In setting (2), our method achieved 68% overall accuracy, while the base model reaches 38% (blue square vs. red hollow square), $p=$3.05e-05 in the McNemar test at the 0.05 level. In setting (3) we observe a 80% overall accuracy for our method vs. 52% for the base model (blue triangle vs. red hollow triangle). This improvement is also statistically significant, with $p=$1.22e-04 in the McNemar test at the 0.05 level. That is, analogy retrieval boosts model performance across all settings.
Verifier feedback. We now measure the impact of retries following feedback. As shown in Figure 3, allowing retries consistently improves results across all difficulty levels. For our analogy-based method, retries (blue square) yields an average gain of additional 20% over no retries (blue circle), with gains ranging from 10% to 30% at different levels. The improvement is even more pronounced for the base model (red hollow square vs. red hollow circle), where verifier feedback results in an average gain of 26%, ranging from 10% to 40% per level.
@@
Effect of $m$ and $n$ parameters. Figure 4 shows the average number of retries (top) and runs (bottom) per problem by difficulty level. Dashed line represents maximum allowed, but the method stops early if it reaches a valid proof. Our method consistently requires fewer retries and runs than the base model across all levels. On average, it uses 4.7 retries and 1.6 runs per problem, compared to 9.02 retries and 2.14 runs for the base model.
Proof accuracy given correct answers. We consider a numeric answer to be correct if it is correct in at least one retry across runs. By this definition, our full pipeline solved 94% of problems (47). Among these, it generated correct proofs for 40 problems (85.1% success rate, compared to 80% across all problems). In contrast, the base model (multiple runs, retries) produced 44 correct answers (88%), with only 26 correct proofs (59% success rate, compared to 52% on all problems).
Interestingly, while the base model often finds the right answer, it fails to support it with a correct proof. We conjecture that the model often has some notion of the correct proof but struggles to precisely phrase it. By providing relevant proofs from analogous problems, our method helps the model close this gap, which is crucial for trustworthy systems. Notably, our method also improves the overall accuracy in finding the correct numerical answer (94% vs. 88%).
Stability of results. Our evaluation was conducted on a set of 50 problems (10 per difficulty level 1-5). To assess the stability of our results, we sample an additional 10 problems per level and rerun our method. Due to its high cost, we do not repeat this for the base-model baseline. [^2] The baseline includes the entire theorem dictionary in the prompt, leading to high API call costs. We observe that overall accuracy varies by only 4% per level on average, and conclude the results are stable. See Appendix 9 for per-level differences.
@@ 5.4 Error Analysis
As expected, our method yields fewer errors than the base model across all tiers and levels, except for tier-2 at level 4. Tier-1 syntax errors are the most prevalent. These are followed by tier-3 errors, where the proof fails to conclude the goal. Tier-2 errors – calling a theorem without satisfying one of the premises, are the least common. This trend holds for the base model as well. As noted in Section 3.4, tier-1 errors are frequent but easily resolved by the LLM after feedback.
@@ 6 Related Work
Neuro-symbolic Methods. Neuro-symbolic approaches combine LLMs with symbolic tools. This strategy has been shown to be effective for some structured tasks Zhou et al. (2024); Trivedi et al. (2024); Zhang et al. (2024); Yuan et al. (2024). Auto-formalization methods Chen and Szegedy (2024); Fischer and Kaliszyk (2024) such as LINC Olausson et al. (2023) translate natural language into first-order logic using an LLM, followed by a symbolic theorem prover. We note that the focus of these works is on logical reasoning, and they are not readily equipped to handle the mathematical or algebraic elements within proofs.
@@ 7 Conclusions and Future Work
    On the other hand, we note that a recent work Xu et al. (2024) has shown that multimodal LLMs often struggle with the visual aspects of math problems, often performing better without the images.
@@ Appendix C Verifier Implementation Details
  - Error: You tried to use theorem: isosceles_triangle_judgment_line_equal(1,BOD);
Polygon(BOD)&Equal(LengthOfLine(BO),LengthOfLine(BD)); [’IsoscelesTriangle(BOD)’]
Missing premise: Polygon for triangle BOD is missing
Details: The construction data did not define this polygon.
@@
Collinear Points: ACD, AOB
Polygons: ACO, ADO, AOC, AOD, CDO, COD
- Theorems related to the goal: None found that constrain this goal
- Solver constraints directly related to this goal: None found
@@ Appendix D Stability of Results
Figure 9: Accuracy remains stable when increasing the sample size from 50 to 100 (10 additional problems per level), with an average variation of only 4% per level. Per-level differences range from 0% (levels 2, 3, and 5) to 10% (levels 1 and 4). We conclude our results are stable.
Figure 9 shows the performance of our method on both the original 50 samples and the extended set of 100 samples, broken down by level. As shown, performance remains unchanged in levels 2, 3, and 5, while levels 1 and 4 show a 10% difference. Overall, the average difference across levels is 4%. We conclude our results are stable.

### v3

@@ 1 Introduction
Figure 2: An example problem from FormalGeo-7k (level 3). Left: Problem inputs, including a natural language description and formal representations – construction (entities and relations), extended construction (inferred based on extension rules), conditions, and goal. The input also includes a numeric answer (or expression) and a formal proof, composed of steps, each invoking a theorem from the dictionary. Right: The Theorem Dictionary. Theorems include variation id, arguments, premises and conclusions. The full dataset also includes diagrams, but we do not use them here. A recent work Zhang et al. (2024) has shown that multimodal LLMs often struggle with the visual aspects of math problems, often performing better without the images.
@@
Mathematical proofs demand rigorous logical deduction, symbolic manipulation, and an understanding of abstract concepts that go beyond statistical correlations in language. The requirement for absolute truth and the absence of ambiguity in mathematical reasoning present a significant challenge for models trained to generate plausible text rather than formally valid inferences Singh et al. (2024); Pan et al. (2025). In addition, the often lengthy nature of proofs necessitates a level of sustained logical coherence and hierarchical reasoning that is hard for current LLMs.
Enabling LLMs to generate rigorous and verifiable proofs can dramatically boost LLMs’ reliability, accuracy and consistency, unlocking applications in mathematics, science and education Welleck et al. (2022); Gupta et al. (2025); Kumar et al. (2023), as well as many safety-critical and security-critical tasks.
@@
As a proof of concept, we focus on Euclidean geometry – a domain that is symbolic, verifiable, and rich in structural analogies. Note that our primary objective in this paper is to evaluate whether our symbolic augmentations can improve the proof-generation capabilities of general-purpose LLMs. As such, while geometry is the domain we test our ideas in, our focus is not on competing with state-of-the-art, specialized geometry solvers. Rather, we wish to quantify the gains enabled by our method.
Our main contributions are:
@@ 4 Experimental Setup
Base model. We use OpenAI’s o1 model with the default parameters as our base model. o1 is a state-of-the-art LLM known for its capabilities in mathematical reasoning, proof generation, and complex problem-solving Jaech et al. (2024). We experimented with several models, including GPT-4, GPT-4o-mini, and GPT-4o Achiam et al. (2023); Hurst et al. (2024), and observed that these models tend to produce proofs with a significant number of syntax errors. o1 consistently outperformed them, making it the strongest baseline for our task. We leave the evaluation of additional base models for future work.
@@ 5 Results
In setting (1) our method achieved an overall accuracy of 48%, substantially outperforming the baseline’s 10% (blue circle vs. red hollow circle), $p=$3.81e-06 in the McNemar test at the 0.05 level. In setting (2), our method achieved 68% overall accuracy, while the base model reaches 38% (blue square vs. red hollow square), $p=$6.10e-05 in the McNemar test at the 0.05 level. In setting (3) we observe a 80% overall accuracy for our method vs. 52% for the base model (blue triangle vs. red hollow triangle). This improvement is also statistically significant, with $p=$1.22e-04 in the McNemar test at the 0.05 level. That is, analogy retrieval boosts model performance across all settings.
Verifier feedback. We now measure the impact of retries following feedback. As shown in Figure 3, allowing retries consistently improves results across all difficulty levels. For our analogy-based method, retries (blue square) yields an average gain of additional 20% over no retries (blue circle), with gains ranging from 10% to 30% at different levels. The improvement is even more pronounced for the base model (red hollow square vs. red hollow circle), where verifier feedback results in an average gain of 28%, ranging from 10% to 40% per level.
@@
Effect of $m$ and $n$ parameters. Figure 4 shows the average number of retries (top) and runs (bottom) per problem by difficulty level. Dashed line represents maximum allowed, but the method stops early if it reaches a valid proof. Our method consistently requires fewer retries and runs than the base model across all levels. On average, it uses 4.6 retries and 1.58 runs per problem, compared to 8.92 retries and 2.12 runs for the base model.
Proof accuracy given correct answers. We consider a numeric answer to be correct if it is correct in at least one retry across runs. By this definition, our full pipeline solved 100% of problems (50). Among these, it generated correct proofs for 40 problems (80% success rate). In contrast, the base model (multiple runs, retries) produced 45 correct answers (90%), with only 26 correct proofs (57.7% success rate, compared to 52% on all problems).
Interestingly, while the base model often finds the right answer, it fails to support it with a correct proof. We conjecture that the model often has some notion of the correct proof but struggles to precisely phrase it. By providing relevant proofs from analogous problems, our method helps the model close this gap, which is crucial for trustworthy systems. Notably, our method also improves the overall accuracy in finding the correct numerical answer (100% vs. 90%).
Stability of results. Our evaluation was conducted on a set of 50 problems (10 per difficulty level 1-5). To assess the stability of our results, we sample an additional 10 problems per level and rerun our method. Due to its high cost, we do not repeat this for the base-model baseline. [^2] The baseline includes the entire theorem dictionary in the prompt, leading to high API call costs. We observe that overall accuracy varies by only 3% per level on average, and conclude the results are stable. See Appendix 9 for per-level differences.
@@ 5.4 Error Analysis
As expected, our method yields fewer errors than the base model across all tiers and levels. Tier-1 syntax errors are the most prevalent. These are followed by tier-3, where the proof fails to conclude the goal. Tier-2 – calling a theorem without satisfying one of the premises, are the least common. This trend holds for the base model as well. As noted in Section 3.4, tier-1 errors are frequent but easily resolved by the LLM after feedback.
@@ 6 Related Work
Neuro-symbolic Methods. Neuro-symbolic approaches combine LLMs with symbolic tools. This strategy has been shown to be effective for some structured tasks Szeider (2024); Régin et al. (2024); Mittal et al. (2024); Wu and Mitra (2024). Auto-formalization methods Song et al. (2024); Alhessi et al. (2025) such as LINC Olausson et al. (2023) translate natural language into first-order logic using an LLM, followed by a symbolic theorem prover. We note that the focus of these works is on logical reasoning, and they are not readily equipped to handle the mathematical or algebraic elements within proofs.
@@ 7 Conclusions and Future Work
    On the other hand, we note that a recent work Zhang et al. (2024) has shown that multimodal LLMs often struggle with the visual aspects of math problems, often performing better without the images.
@@ Appendix C Verifier Implementation Details
  - Error: You tried to use theorem: right_triangle_judgment_angle(1,BCD);
Polygon(BCD)&Equal(MeasureOfAngle(BCD),90);[’RightTriangle(BCD)’]
Missing premise: Angle measure 90° for triangle BCD is not established in the premise.
Details: Premise provided: Polygon(BCD)&Equal(MeasureOfAngle(BCD),90)
@@
Perpendicular Lines: AD, BD
Collinear Points: ADC
Polygons: ABC, ABD, ACB, ADB, BCD, BDC
- Theorems related to the goal:
Step 1 - right_triangle_judgment_angle(1, BCD): RightTriangle(BCD)
- Solver constraints directly related to this goal:
|AB| = y
@@ Appendix D Stability of Results
Figure 9: Accuracy remains stable when increasing the sample size from 50 to 100 (10 additional problems per level), with an average variation of only 3% per level. Per-level differences range from 0% (levels 2, 3, and 5) to 5% (level 1), and 10% (level 4). We conclude our results are stable.
Figure 9 shows the performance of our method on both the original 50 samples and the extended set of 100 samples, broken down by level. As shown, performance remains unchanged in levels 2, 3, and 5, while levels 1 show 5% difference, and level 4 4 show a 10% difference. Overall, the average difference across levels is 3%. We conclude our results are stable.